# Is Stephen Curry really a guard? New perspective on players typology using functional data analysis

### v1

@@ 4.2 Clustering
The $k$-medoids algorithm was applied to the coefficients plotted in Figure 5, grouping players into five clusters using an euclidean distance and different weighting schemes for the components. The choice of five clusters aligns with the number of players on the court for each team. Detailed clustering results, including player names and decomposition in eigenfunctions basis, are provided in Appendix A. Figure 6 presents the medoids for each cluster when the components are weighted equally. The medoid, as the central observation in a cluster, represents the most representative player within that cluster. The decompositions of these representative players are given in Figure 9 in the Appendix. The clustering results can be interpreted based on shooting tendencies. Cluster $1$ groups high-volume shooters. Cluster $3$ includes players who primarily shoot from the center axis of the basket, while players in Cluster $4$ predominantly score from the corners. Clusters $2$ and $5$ are similar, comprising players who primarily score in the paint but also attempt shots from beyond the three-point line, albeit with limited success.
@@
Figure 6: Most representative player of each cluster according to the $k$-medoids algorithm using an euclidean distance with equal weights for each components.
@@
Figure 7 presents the cluster medoids for the results obtained with components weighted using the percentage of variance they explain. Since the first component accounts for approximately $70\%$ of the total variance, the clusters are primarily driven by the first score $c_{1}$, reflecting the overall shooting frequency of the players. The clusters span a spectrum of shooting behaviors, from players who predominantly score in the paint (Cluster $3$) to those who are prolific three-point shooters (Cluster $4$). This effect is also evident in the decomposition shown in Figure 10 in the Appendix. The clusters can be ranked relative to the average player’s shooting tendencies. Compared to the average player, Cluster $3$ players shoot more frequently in the paint and attempt fewer three-pointers. Players in Cluster $1$ shoot slightly less than the average player, while those in Cluster $5$ align with the average. Cluster $2$ comprises players who shoot slightly more than average, and Cluster $4$ players favor three-pointers significantly more.
@@
Figure 7: Most representative player from each cluster according to the $k$-medoids algorithm using an euclidean distance with weights defined using the percentage of variance explained by each component.
@@
Figure 8 presents the differences between the NBA-defined typology and the two clustering approaches we present. The NBA clusters are based on the usual player classification (00: center, $1$: forward, $2$: forward-center, $3$: forward-guard, $4$: guard). Overall, there is minimal agreement between the NBA clusters and the $k$-medoids clusters, whether weights are equal or based on the percentage of variance explained. The only exception is the NBA-defined ‘center’ category, which aligns with the cluster $2$ in the $k$-medoids results when weighted by the variance explained. Furthermore, there is little agreement between the two $k$-medoids clustering methods themselves.
@@
Figure 8: Agreement between the two clustering taxonomies (using equal weights and variance weights) and the conventional NBA position typology. NBA clusters are based on the usual typology (00: center, $1$: forward, $2$: forward-center, $3$: forward-guard, $4$: guard).
@@ 5 Discussion
We envision a number of potential uses of both our functional shot-chart representation and the subsequent clustering of shooting patterns. Firstly, the functional representation of a players shot-chart provides a comprehensive characterization of a player’s shooting behavior, i.e., the shot chart is not reduced to one or more scalar summaries. This representation could be used by coaches or front-office staff to monitor changes in a player’s shooting patterns between or within seasons, to identify player-specific tendencies which inform specific training regimes, and also to identify players whose characteristics match the needs of the team during scouting and recruitment. Because the charts can naturally be presented in a way that is interpretable to basketball fans who may not be familiar with the underlying mathematics and statistics, the representation can be used to tell stories to the general public about player tendencies, and enable historical comparisons between players, enriching the narrative around the sport. While we used shots made and shots missed, other metrics defined over the offensive half-court (e.g., shooting efficiency or defensive shot charts, Franks et al., (2015)), could be represented in the same way for the aforementioned purposes.
@@
There are also a number of possible methodological extensions to the modeling framework. One of the most exiting avenues is to combine the shot-chart representations with other high-throughput data sources that are now routinely collected within the NBA. For example, temporal biomechanical data are now being measured by motion-tracking systems such as Hawk-Eye, and the possibility of a multi-modal representation that reflects a players tactical and physical tendencies – and the interaction between them – could provide more detailed and specific analyses. Our modeling framework could also be applied to positional data in other sports, for example, to characterize a soccer player’s “natural position” or a boxer’s “fighting style”.
Appendix A Clustering results
@@
Cluster 3
Cluster 4
@@
Cluster 2
@@
Cluster 3
Cluster 4
@@
Cluster 5

### v2

@@ 4.2 Clustering
The $k$-medoids algorithm was applied to the coefficients plotted in Figure 5, grouping players into five clusters using a Euclidean distance and different weighting schemes for the components. The choice of five clusters aligns with the number of players on the court for each team. Detailed clustering results, including player names and decomposition in eigenfunctions basis, are provided in Appendix C. Figure 6 presents the medoids for each cluster when the components are weighted equally. The medoid, as the central observation in a cluster, represents the most representative player within that cluster. The decompositions of these representative players are given in Figure 15 in the Appendix. The clustering results can be interpreted based on shooting tendencies. Cluster $1$ groups high-volume perimeter shooters. Cluster $2$ includes players who primarily shoot from the center axis of the basket (represented by ), while players in Cluster $3$ predominantly score from the corners. Clusters $4$ and $5$ are similar, comprising players who primarily score in the paint but also attempt shots from beyond the three-point line, albeit with limited success.
@@
Figure 6: Most representative player of each cluster according to the $k$-medoids algorithm using a Euclidean distance with equal weights for each component.
@@
Figure 7 presents the cluster medoids for the results obtained with components weighted using the percentage of variance they explain. Since the first component accounts for approximately $70\%$ of the total variance, the clusters are primarily driven by the first score $c_{1}$, reflecting the overall shooting frequency of the players. The clusters span a spectrum of shooting behaviors, from players who predominantly score in the paint (Cluster $3$) to those who are prolific three-point shooters (Cluster $4$). This effect is also evident in the decomposition shown in Figure 16 in the Appendix. The clusters can be ranked relative to the average player’s shooting tendencies. Compared to the average player, Cluster $1$ players shoot more frequently in the paint and attempt fewer three-pointers. Players in Cluster $2$ shoot slightly less than the average player, while those in Cluster $3$ align with the average. Cluster $4$ comprises players who shoot slightly more than average, and Cluster $5$ players favor three-pointers significantly more.
@@
Figure 7: Most representative player from each cluster according to the $k$-medoids algorithm using a Euclidean distance with weights defined using the percentage of variance explained by each component.
@@
Figure 8 presents the confusion matrices between the NBA-defined typology and the two clustering approaches we present. The NBA clusters are based on the usual player classification. Overall, there is minimal agreement between the NBA clusters and the $k$-medoids clusters, whether weights are equal or based on the percentage of variance explained. The Adjusted Rand Index (ARI, hubertComparingPartitions1985) between the NBA clusters and the $k$-medoids clusters with equal weights is $0.05$ and between the NBA clusters and the $k$-medoids clusters with variance weights is $0.04$. This is expected as NBA clusters are declarative and not data-based. The only exception is the NBA-defined ‘center’ category, which aligns with the cluster $1$ in the $k$-medoids results when weighted by the variance explained. Cluster $1$ of the weighted $k$-medoids contains players that shoot mostly under the basket, which is coherent with the usual scoring position of centers. Furthermore, there is little agreement between the two $k$-medoids clustering methods themselves. The ARI is equal to $0.1$. The Silhouette coefficient (rousseeuwSilhouettesGraphicalAid1987) for the $k$-medoids with equal weights is equal to $0.04$, while the Silhouette coefficient for the $k$-medoids with variance weights is equal to $0.4$. We also computed the Silhouette coefficient of the NBA clustering using the distance matrix used for the $k$-medoids with equal weights (0.05-0.05) and using the distance matrix used for the $k$-medoids with variance weights (0.1-0.1).
@@
Figure 8: Agreement between the two clustering taxonomies (using equal weights and variance weights) and the conventional NBA position typology.
@@ 5 Discussion
We envision a number of potential uses of both our functional shot-chart representation and the subsequent clustering of shooting patterns. Firstly, the functional representation of a player’s shot chart provides a comprehensive characterization of a player’s shooting behavior, i.e., the shot chart is not reduced to one or more scalar summaries. This representation could be used by coaches or front-office staff to monitor changes in a player’s shooting patterns between or within seasons, to identify player-specific tendencies which inform specific training regimes, and also to identify players whose characteristics match the needs of the team during scouting and recruitment. Because the charts can naturally be presented in a way that is interpretable to basketball fans who may not be familiar with the underlying mathematics and statistics, the representation can be used to tell stories to the general public about player tendencies, and enable historical comparisons between players, enriching the narrative around the sport. While we used shots made and shots missed, other metrics defined over the offensive half-court (e.g., shooting efficiency or defensive shot charts, franksCharacterizingSpatialStructure2015), could be represented in the same way for the aforementioned purposes.
@@
There are also a number of possible methodological extensions to the modeling framework. One of the most exciting avenues is to combine shot-chart representations with other high-throughput data sources that are now routinely collected within the NBA. For example, temporal biomechanical data are now being measured by motion-tracking systems such as Hawk-Eye (HawkeyeInnovation2025), and the possibility of a multi-modal representation that reflects a player’s tactical and physical tendencies – and the interaction between them – could provide more detailed and specific analyses. Our modeling framework could also be applied to positional data in other sports, for example, to characterize a soccer player’s “natural position” or a boxer’s “fighting style”.
Finally, some limitations of the proposed modelling framework are as follows. We employed standard FPCA which assumes that observations are realizations from an unconstrained space of functions. However, densities are constrained because they must be non-negative and integrate to $1$. Reconstructions from standard FPCA do not necessarily satisfy the non-negativity constraint so specialized methods have been developed (petersenFunctionalDataAnalysis2016). However, these are for the univariate case. Rather than develop or extend these methodological approaches to the bivariate case, we apply the more tractable, standard FPCA approach while acknowledging its limitations. If required, reconstructions that lie outside the space of densities can be projected back into the density space. Secondly, this work focused on characterising the spatial patterns of made or missed shots. The normalization in the density estimation step, performed separately for the made and missed shots, inherently removes information about total shot volume or shooting efficiency (e.g., percentage chance of success). Future work could include an overall measure of shot volume or shooting efficiency to improve the characterization of players’ behaviors.
Appendix A Court description
Here, we present a description of the spatial regions of the basketball court referenced throughout the main text. These regions correspond to functionally important areas of play. Figure 9 presents the important locations on the offensive half-court. The basket is located at the center of the baseline, with the low-post regions surrounding it on both the left and right sides. Above these areas, the high-post zones are located near the free-throw line. The three-point line is the line that separates the two-point area from the three-point area, i.e, any shot made beyond this line counts as three point. The zone located between the free-throw line and the three-point line is commonly referred to as the top of the key. The corners of the court are divided into short-corner zones, that are closer to the basket, and corner three-point zones along the baseline. Finally, the wing areas are located on the side of the court between the sidelines and the extension of the free-throw line.
Figure 9: Description of the basketball court.
Appendix B Bootstrap results
As a sensitivity check, we performed the following exercise:
  • Construct $5$ bootstrap samples.
  • For each bootstrap sample, compute the MFPCA.
  • Display the mean and the first 44 MFPCs from each bootstrap sample (Figures 10 to 14).
The shapes of the resulting mean functions and the first two principal components are very consistent. There is slight sample-to-sample variability for the third component. This result is intuitive. The mean and leading components, explaining the most variability, represent the most common modes of variation in the data and hence are more stable from sample to sample. Overall, this simple sensitivity check indicates that the prominent modes of variation are stable.
Figure 10: Bootstrap: First sample results.
Figure 11: Bootstrap: Second sample results.
Figure 12: Bootstrap: Third sample results.
Figure 13: Bootstrap: Fourth sample results.
Figure 14: Bootstrap: Fifth sample results.
Appendix C Clustering results
@@
Number of guards: 18; forward-guard: 5; forward: 7; forward-center: 2; center: 1.
@@
Number of guards: 28; forward-guard: 3; forward: 11; forward-center: 7; center: 8.
Cluster 3
Number of guards: 4; forward-guard: 9; forward: 11; forward-center: 0; center: 0.
Cluster 4
@@
Number of guards: 4; forward-guard: 4; forward: 8; forward-center: 3; center: 0.
@@
Number of guards: 14; forward-guard: 3; forward: 6; forward-center: 13; center: 4.
@@
Number of guards: 4; forward-guard: 1; forward: 6; forward-center: 8; center: 11.
Cluster 2
@@
Number of guards: 12; forward-guard: 9; forward: 12; forward-center: 10; center: 1.
Cluster 3
Number of guards: 14; forward-guard: 4; forward: 9; forward-center: 3; center: 0.
Cluster 4
@@
Number of guards: 19; forward-guard: 4; forward: 7; forward-center: 3; center: 1.
Cluster 5
@@
Number of guards: 19; forward-guard: 6; forward: 9; forward-center: 1; center: 0.